OpenStudy (sugarapples):

3 √-32xy^6z^8 need help please

OpenStudy (anonymous):

\[\sqrt[3]{-32xy ^{6}z ^{8}}\] \[\sqrt{4} = 2^(2-2) = 2^1 = 2\]

OpenStudy (anonymous):

Treat the cube root s something you subtract the exponents from. \[\sqrt[3]{9^{5}} = 9^(5-3) = 9^2 = 81\]

OpenStudy (anonymous):

the x has an exponent of 1

OpenStudy (anonymous):

So it would be: -32x^(1-3) y^(6-3) z^(8-3)